# Towards Consistent Object Detection via LiDAR-Camera Synergy

### v1

@@ Abstract
As human-machine interaction continues to evolve, the capacity for environmental perception is becoming increasingly crucial. Integrating the two most common types of sensory data, images, and point clouds, can enhance detection accuracy. However, currently, no model exists that can simultaneously detect an object’s position in both point clouds and images and ascertain their corresponding relationship. This information is invaluable for human-machine interactions, offering new possibilities for their enhancement. In light of this, this paper introduces an end-to-end Consistency Object Detection (COD) algorithm framework that requires only a single forward inference to simultaneously obtain an object’s position in both point clouds and images and establish their correlation. Furthermore, to assess the accuracy of the object correlation between point clouds and images, this paper proposes a new evaluation metric, Consistency Precision (CP). To verify the effectiveness of the proposed framework, an extensive set of experiments has been conducted on the KITTI and DAIR-V2X datasets. The study also explored how the proposed consistency detection method performs on images when the calibration parameters between images and point clouds are disturbed, compared to existing post-processing methods. The experimental results demonstrate that the proposed method exhibits excellent detection performance and robustness, achieving end-to-end consistency detection. The source code will be made publicly available at https://github.com/xifen523/COD.
@@ I Introduction
Human-machine interaction necessitates the perception of the surrounding environment, and object detection is one of the most commonly employed methods of perception. The task of object detection [1, 2], a classic endeavor in computer vision, has historically garnered extensive interest from numerous researchers. Depending on the data type, object detection can be categorized into 2D object detection [3] based on images and 3D object detection [4] based on point clouds, each with broad applications. For certain specialized tasks aimed at achieving higher detection accuracy or robustness, there is also 3D object detection [5] that integrates both images and point clouds.
Although existing detectors [6, 7] are powerful, they struggle to establish the correspondence of targets across multiple modalities. Even when the input data consists of images and point clouds, the final detection result is the object’s bounding box in the 3D point cloud, without a corresponding bounding box in the 2D image. To simultaneously obtain the target’s location in the 3D point cloud and in the image, further post-processing steps are required. There are typically two implementation methods. The first method utilizes the calibration matrix between the point cloud and the camera to calculate the positions of the eight corners of the 3D bounding box in the image and then infers the largest bounding box based on these eight corners [8]. The second method involves using a 2D image detector to match the 2D detection results of the image with the 3D detection results, thereby obtaining both 2D and 3D detection outcomes for the object.
@@ II-A Visual Object Detection
Visual object detection has been developed for a long time, and traditional object detection techniques use manual feature extraction with poor robustness and high computational time complexity, which limits the improvement of object detection algorithms. With the continuous emergence of deep learning methods, Fast R-CNN [12], Faster R-CNN [13], and their related improved algorithms [14, 15] have achieved excellent results as two-stage object detection approach effect. The single-stage object detector predicts the object category and location directly from the image, which has the advantages of fast speed and fewer computational resources.
SSD [16] is the first single-stage object detection method that achieves similar detection accuracy to Faster R-CNN in real-time performance. Subsequently, a series of YOLO-based methods [17, 18, 19] are proposed to further optimize and improve the performance of single-stage object detection methods.
Recent research work has also proposed the use of end-to-end prediction methods to directly predict the bounding box of each object. Among them, DETR [9] is the most representative end-to-end prediction method, which directly predicts the location as well as the category of an object by decoding a learnable query. Since then, a series of improved DETR-based methods [20, 21, 22] keeps emerging. End-to-end prediction is gradually becoming the dominant paradigm in the field of object detection due to the removal of the complex post-processing process.
@@ II-B Point Cloud Object Detection
In addition, the point-based processing is closer to the original LiDAR data and does not lose the original geometric information due to quantization errors. PointNet++ [27] applies PointNet [28] in a hierarchical recursive manner to improve the detection performance by adaptively capturing the structure and fine-grained features of the point cloud. fine-grained features to improve detection performance. Point RCNN [29] proposes a point-based two-stage detection framework that extends the classical 2D object detection framework, Faster R-CNN, to 3D object detection. 3DSSD [30] proposes a lightweight framework that achieves a balance between detection accuracy and speed.
@@ II-C Fusion-based Object Detection
Mid-term fusion, which occurs after data preprocessing has been completed and before final detection results are generated, is currently the method with the most potential for development. TransFusion [33] defines object queries in 3D space and fuses image features into these proposal boxes. MMF [34] utilizes multiple related subtasks (e.g., ground estimation, depth completion, 2D object detection, etc.) to enhance cross-modal fusion learning. DeepFusion [35] proposes InverseAug inverse rotation and other geometrically relevant data enhancements to achieve precise geometric alignment between LiDAR points and image pixels. SFD [36] enhances LiDAR-generated sparse point clouds using a dense pseudo-point cloud generated by depth complementation, which fuses 3D RoI features from different point clouds into a 3D mesh. 4D-Net [36] places the fusion module in the point cloud feature extractor to allow the point cloud features to be dynamically focused on image features.
@@ IV-A2 Metrics
IV-A3 Implementation Details
In this study, the 3D object detection networks—CenterPoint [38], PillarNet [39], SECOND [24], and PointPillar [25]—were trained using the original models from the open-source OpenPCDet [8] repository, with the standard configuration files and trained over $80$ epochs. For the 2D detection, the RT-DETR [3] code from the open-source YOLOV8 [40] project was employed. The consistency detection architecture involved integrating the RT-DETR code from YOLOV8 into OpenPCDet. In all of our experiments, whether conducted on the KITTI or DAIR-V2X datasets, we configured the point cloud and image consistency detection algorithm training parameters as follows: the network input image size scaling was set to $640{\times}640$, and the range of the point cloud was from the current to the $70.4$ meter position, $40$ meters from left to right, and from ${-}3$ to $1$ meter from top to bottom.
All experiments were conducted on a system running Ubuntu 20.04, equipped with two GeForce RTX 3090 GPUs. The experiments ran for $80$ epochs using the “Adam+onecycle” optimizer. The learning rate was set at $0.003$, weight decay at $0.01$, and momentum at $0.9$. Additionally, the batch size was configured to $4$, with all other settings remaining at their default values. Unless specified otherwise, the noise referred to in the experiments pertains to the addition of noise to the extrinsic matrix between the point cloud and image data. Specifically, this noise consists of random rotational disturbances around the X, Y, and Z axes with a mean of 00 and a variance of $0.01$ degrees, accompanied by random translational disturbances along the X, Y, and Z axes with a mean of 00 and a variance of $0.0005$ meters. Additionally, the DAIR-V2X dataset can be converted into the KITTI dataset format using its own scripts, allowing it to maintain the same settings and utilize the same tools as the KITTI dataset. It is worth noting that the Bbox mAP of the evaluated 3D detectors [38, 25, 39, 24] in this paper is derived using post-processing methods from the detected 3D bounding boxes.
@@ IV-B2 Consistency Precision Analysis
We have evaluated the accuracy of target correspondence within the consistency method between images and point clouds and compared it to traditional match-based methods for determining correspondences. These traditional methods first use separate 3D and 2D detectors to identify targets and then employ the calibration matrix between point clouds and images to match 3D detection boxes with 2D detection boxes. In this study, the 2D detector employed was RT-DETR [3] from YOLOV8 [40].
@@ IV-C1 Noise Resistance of Different Methods
Figure 3: (a) and (b) depict the visualization of object bounding boxes on images using the postprocessing method and the consistency detection method, under noisy and noise-free settings, respectively.
@@ IV-C2 Different Noise Effects
IV-D Visualisation Results
We have visually compared the traditional method of calculating image bounding boxes (Bbox) from 3D Bbox data with our proposed method, as presented in Figure 3. The comparison also examines the performance of both methods under conditions with and without noise. As depicted in Figure 3 (b), despite the presence of noise, the detection boxes of the COD method continue to align closely with the objects, thus reinforcing the viability of our COD approach in noisy environments. Additionally, contrasting Figure 3 (a) with Figure 3 (b) reveals that while post-processing-based methods experience considerable displacement of bounding boxes following noise introduction, our COD method consistently maintains tightly fitted detection boxes around the targets, irrespective of noise, thereby demonstrating superior robustness.

### v2

@@ Abstract
As human-machine interaction continues to evolve, the capacity for environmental perception is becoming increasingly crucial. Integrating the two most common types of sensory data, images, and point clouds, can enhance detection accuracy. Currently, there is no existing model capable of detecting an object’s position in both point clouds and images while also determining their corresponding relationship. This information is invaluable for human-machine interactions, offering new possibilities for their enhancement. In light of this, this paper introduces an end-to-end Consistency Object Detection (COD) algorithm framework that requires only a single forward inference to simultaneously obtain an object’s position in both point clouds and images and establish their correlation. Furthermore, to assess the accuracy of the object correlation between point clouds and images, this paper proposes a new evaluation metric, Consistency Precision (CP). To verify the effectiveness of the proposed framework, an extensive set of experiments has been conducted on the KITTI and DAIR-V2X datasets. The study also explored how the proposed consistency detection method performs on images when the calibration parameters between images and point clouds are disturbed, compared to existing post-processing methods. The experimental results demonstrate that the proposed method exhibits excellent detection performance and robustness, achieving end-to-end consistency detection. The source code will be made publicly available at https://github.com/xifen523/COD.
@@ I Introduction
Human-machine interaction necessitates the perception of the surrounding environment, and object detection is one of the most commonly employed methods of perception. Depending on the data type, object detection can be categorized into 2D object detection [1] based on images and 3D object detection [2] based on point clouds, each with broad applications. For certain specialized tasks aimed at achieving higher detection accuracy or robustness, there is also 3D object detection [3] that integrates both images and point clouds.
Although existing detectors [4, 5] are powerful, they struggle to establish the correspondence of targets across multiple modalities. Even when the input data consists of images and point clouds, the final detection result is the object’s bounding box in the 3D point cloud, without a corresponding bounding box in the 2D image. To simultaneously obtain the target’s location in the 3D point cloud and in the image, further post-processing steps are required. There are typically two implementation methods. The first method utilizes the calibration matrix between the point cloud and the camera to calculate the positions of the eight corners of the 3D bounding box in the image and then infers the largest bounding box based on these eight corners. The second method involves using a 2D image detector to match the 2D detection results of the image with the 3D detection results, thereby obtaining both 2D and 3D detection outcomes for the object.
@@ II-A Visual Object Detection
Visual object detection has evolved significantly. Traditional techniques with manual feature extraction suffered from poor robustness and high computational complexity. Deep learning methods, such as Fast R-CNN [9], Faster R-CNN [10], and their improved versions [11, 12], have achieved excellent results as two-stage detection approaches. Single-stage detectors predict object categories and locations directly from the image, offering speed and lower computational demands.
SSD [13] was the first single-stage method to achieve real-time performance with accuracy comparable to Faster R-CNN. Following this, YOLO-based methods [14, 15] have further optimized single-stage detection.
Recent work has introduced end-to-end methods for directly predicting object bounding boxes. DETR [6] is a key example, using learnable queries to predict object locations and categories. Improved DETR-based methods [16, 17, 18] continue to emerge, making end-to-end prediction a dominant paradigm by eliminating complex post-processing.
@@ II-B Point Cloud Object Detection
In addition, the point-based processing is closer to the original LiDAR data and does not lose the original geometric information due to quantization errors. PointNet++ [23] applies PointNet in a hierarchical recursive manner to improve the detection performance by adaptively capturing the structure and fine-grained features of the point cloud. fine-grained features to improve detection performance. Point RCNN [24] proposes a point-based two-stage detection framework that extends the classical 2D object detection framework, Faster R-CNN, to 3D object detection. 3DSSD [25] proposes a lightweight framework that achieves a balance between detection accuracy and speed.
@@ II-C Fusion-based Object Detection
Mid-term fusion, which occurs after data preprocessing has been completed and before final detection results are generated, is currently the method with the most potential for development. TransFusion [28] defines object queries in 3D space and fuses image features into these proposal boxes. DeepFusion [29] proposes InverseAug inverse rotation and other geometrically relevant data enhancements to achieve precise geometric alignment between LiDAR points and image pixels. SFD [30] enhances LiDAR-generated sparse point clouds using a dense pseudo-point cloud generated by depth complementation, which fuses 3D RoI features from different point clouds into a 3D mesh. 4D-Net [30] places the fusion module in the point cloud feature extractor to allow the point cloud features to be dynamically focused on image features.
@@ IV-B2 Consistency Precision Analysis
We have evaluated the accuracy of target correspondence using the consistency method between images and point clouds and compared it to traditional match-based methods for determining correspondences. These traditional methods first use separate 3D and 2D detectors to identify targets and then employ the calibration matrix between point clouds and images to match 3D detection boxes with 2D detection boxes.